OpenStudy (cedric):

Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

jimthompson5910 (jim_thompson5910):

Hint: Add the equations together to get x - y = 16 2x + y = 2 ---------- 3x + 0y = 18 So 3x = 18

OpenStudy (anonymous):

i think it's like this: to find the value of x you must first get rid of the y, so make whatever number is with y in both equations equal each other and adding them would get rid of them, since one would be positive and one would be negative